Company A's common stock is selling for $75 per share, and currently pays a dividend of$4.75 per share. The share's earnings and dividends are expected to grow at 5% into the foreseeable future. Based on this information, the cost of internal equity financing is:

Group of answer choices

13.20%

11.71%

14.45%

12.71%
Answer #1

Answer:

D0 = $4.75

Growth rate, G = 5%

D1 = D0 *(1+g)
D1 = $4.75 * (1+0.05)
D1 = $4.9875

Cost of Internal Equity Financing = D1 / P0 + g
Cost of Internal Equity Financing = $4.9875 / $75 + 0.05
Cost of Internal Equity Financing = 0.0665 + 0.05
Cost of Internal Equity Financing = 0.1165 or 11.65%